two plastic projectiles have the same mass M and speed V when they hit a glass window pane at a right angle. The interaction between both projectiles and the glass pane takes time T. Projectile A bounces off of the window (in the opposite direction) with the same speed V it had at the beginning. Projectile B sticks to the window pane and stops. Describe the likely damage to the glass pane by selecting all the true statements below.

Projectiles A & B have the same likelihood of breaking the glass since they have the same initial momentum
Projectile B has the greater likelihood of breaking the glass since its final momentum is smaller
Projectile A has the greater likelihood of breaking the glass since its momentum change is larger
Projectile A has the greater likelihood of breaking the glass since its momentum change is smaller
Answer #1

for A: change in momentum = final momentum -initial momentum = -mV-(mv) =-2mv

for B: change in momentum = final momentum -initial momentum = 0-mv=-mv

so, A's change in momentum > B's change in momentum

Force = change in momentum / time

hence A has the greater likelihood of breaking the glass because of larger force

so, C is the correct option